# Progressive Contrastive Learning with Multi-Prototype for Unsupervised Visible-Infrared Person Re-identification

### v1

@@ 3.4 Dynamic Prototype Contrastive Learning
Inspired by MoCo He et al. (2020), we design dynamic prototype contrastive learning in order to preserve the intrinsic diversity in sample features while reducing instability in the simultaneous learning of both common and diverse information. DPCL comprises an online backbone $\phi_{0}$ and a momentum backbone $\phi_{m}$. The momentum backbone mirrors the structure of the online backbone, which is updated by the accumulated weights of the online backbone:
@@
where $\beta$ is a momentum coefficient that controls the update speed of the momentum backbone, which is set to 0.999. $t$ and $t-1$ refer to the current and last iteration, respectively. The momentum backbone ${\phi}_{m}$ update by the moving averaged weights, which are more stable and resistant to sudden fluctuations or noisy updates He et al. (2020).
We randomly select $M$ visible/infrared samples from each cluster, denoted as $X^{v}_{i}$ and $X^{r}_{j}$. We use the momentum backbone $\phi_{m}$ to extract feature representation and store them in visible memory $DM_{RGB}$={$F^{v}_{1}$, $F^{v}_{2}$, $\cdots$, $F^{v}_{N_{v}}$} and infrared memory $DM_{IR}$={$F^{r}_{1}$, $F^{r}_{2}$, $\cdots$, $F^{r}_{N_{r}}$}, as follows:
@@
$dm^{v}_{i}=\begin{cases}\underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\max}\left\|q^{v}-f^{v}_{i}\right\|&\text{if }y_{q}=y_{i}\\ \underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\min}\left\|q^{v}-f^{v}_{i}\right\|&\text{if }y_{q}\neq y_{i}\end{cases},$ (16)
@@
$L_{DPCL}^{v}=-\log\frac{\exp\left(q^{v}\cdot{dm_{+}^{v}}/\tau\right)}{\sum_{i=1}^{N_{v}}\exp\left(q^{v}\cdot{dm_{i}^{v}}/\tau\right)},$ (17)
$L_{DPCL}^{r}=-\log\frac{\exp\left(q^{r}\cdot{dm_{+}^{r}}/\tau\right)}{\sum_{j=1}^{N_{r}}\exp\left(q^{r}\cdot{dm_{j}^{r}}/\tau\right)},$ (18)
$L_{DPCL}=L_{DPCL}^{v}+L_{DPCL}^{r},$ (19)
@@ 3.5 Progressive Contrastive Learning
In the initial training phases, representations are generally of lower quality. Introducing hard samples at this period could be counterproductive, potentially leading the model optimization in an incorrect direction right from the start Zou et al. (2023); Gong et al. (2022). To address this issue, we introduce the Progressive Contrastive Learning, which forms the overall loss function:
${L}_{PCLMP}=\left\{\begin{array}[]{l}{L}_{CPCL},\quad\text{ if }\text{ epoch }\leqslant E_{\text{CPCL }}\\ \lambda{L}_{HPCL}+(1-\lambda){L}_{DPCL},\quad\text{else}\end{array}\right.$ (20)
@@
SVI-ReID DDAG Ye et al. (2020) ECCV’20 54.8 53.0 61.0 68.0 69.4 63.5 68.1 61.8
AGW Ye et al. (2022) TPAMI’21 47.5 47.7 54.2 63.0 70.1 66.4 70.5 65.9
NFS Chen et al. (2021) CVPR’21 56.9 55.5 62.8 69.8 80.5 72.1 78.0 69.8
LbA Park et al. (2021) ICCV’21 55.4 54.1 58.5 66.3 74.2 67.6 72.4 65.5
CAJ Ye et al. (2021) ICCV’21 69.9 66.9 76.3 80.4 85.0 79.1 84.8 77.8
MPANet Wu et al. (2021) CVPR’21 70.6 68.2 76.7 81.0 83.7 80.9 82.8 80.7
DART Yang et al. (2022b) CVPR’22 68.7 66.3 72.5 78.2 83.6 75.7 82.0 73.8
FMCNet Zhang et al. (2022) CVPR’22 66.3 62.5 68.2 74.1 89.1 84.4 88.4 83.9
MAUM Liu et al. (2022) CVPR’22 71.7 68.8 77.0 81.9 87.9 85.1 87.0 84.3
MID Huang et al. (2022) AAAI’22 60.3 59.4 64.9 70.1 87.5 84.9 84.3 81.4
LUPI Alehdaghi et al. (2022) ECCV’22 71.1 67.6 82.4 82.7 88.0 82.7 86.8 81.3
PMT Lu et al. (2023) AAAI’23 67.5 65.0 71.7 76.5 84.8 76.7 84.2 75.1
DEEN Zhang and Wang (2023) CVPR’23 74.7 71.8 80.3 83.3 91.1 85.1 89.5 83.4
SGIEL Feng et al. (2023) CVPR’23 77.1 72.3 82.1 83.0 92.2 86.6 91.1 85.2
PartMix Kim et al. (2023) CVPR’23 77.8 74.6 81.5 84.4 85.7 82.3 84.9 82.5
ProtoHPE Zhang et al. (2023a) MM’23 71.9 70.6 77.8 81.3 88.7 83.7 88.7 82.0
CAL Wu et al. (2023) ICCV’23 74.7 71.7 79.7 83.7 94.5 88.7 93.6 87.6
MUN Yu et al. (2023) ICCV’23 76.2 73.8 79.4 82.1 95.2 87.2 91.9 85.0
SAAI Fang et al. (2023) ICCV’23 75.9 77.0 83.2 88.0 91.1 91.5 92.1 92.0
FDNM Zhang et al. (2024) arXiv’24 77.8 75.1 87.3 89.1 95.5 90.0 94.0 88.7
SSVI-ReID MAUM-50 Liu et al. (2022) CVPR’22 28.8 36.1 - - - - - -
MAUM-100 Liu et al. (2022) CVPR’22 38.5 39.2 - - - - - -
OTLA Wang et al. (2022) ECCV’22 48.2 43.9 47.4 56.8 49.9 41.8 49.6 42.8
TAA Yang et al. (2023b) TIP’23 48.8 42.3 50.1 56.0 62.2 56.0 63.8 56.5
DPIS Shi et al. (2023) ICCV’23 58.4 55.6 63.0 70.0 62.3 53.2 61.5 52.7
USVI-ReID H2H Liang et al. (2021) TIP’21 30.2 29.4 - - 23.8 18.9 - -
OTLA Wang et al. (2022) ECCV’22 29.9 27.1 29.8 38.8 32.9 29.7 32.1 28.6
ADCA Yang et al. (2022a) MM’22 45.5 42.7 50.6 59.1 67.2 64.1 68.5 63.8
NGLR Cheng et al. (2023) MM’23 50.4 47.4 53.5 61.7 85.6 76.7 82.9 75.0
MBCCM He et al. (2023) MM’23 53.1 48.2 55.2 62.0 83.8 77.9 82.8 76.7
CCLNet Chen et al. (2023) MM’23 54.0 50.2 56.7 65.1 69.9 65.5 70.2 66.7
PGM Wu and Ye (2023) CVPR’23 57.3 51.8 56.2 62.7 69.5 65.4 69.9 65.2
GUR* Yang et al. (2023c) ICCV’23 61.0 57.0 64.2 69.5 73.9 70.2 75.0 69.9
MMM Shi et al. (2024) arXiv’24 61.6 57.9 64.4 70.4 89.7 80.5 85.8 77.0
PCLMP Ours 64.4 58.7 69.5 74.4 84.3 80.7 82.7 78.4
PCLMP+MMM Enhanced 65.9 61.8 70.3 74.9 89.6 83.7 87.0 80.9
4 Experiment
We conduct extensive experiments to validate the superiority of our proposed method. First, we provide the detailed experiment setting, which contains datasets, evaluation protocols, and implementation details. Then, we compare our method with many state-of-the-art VI-ReID methods and conduct ablation studies. In addition, to better illustrate our method, we also exhibit further analysis. If not specified, we conduct analysis experiments on SYSU-MM01 in the single-shot & all-search mode.
Table 2: Ablation studies on SYSU-MM01 in all search mode and indoor search mode. “Baseline” means the model trained following PGM  Wu and Ye (2023). Rank-R accuracy(%) and mAP(%) are reported.
Method All Search Indoor Search
@@ 4.1 Experiment Setting
Dataset. We evaluate our method on two common benchmarks in VI-ReID: SYSU-MM01 Wu et al. (2017) and RegDB Nguyen et al. (2017). SYSU-MM01 is a large-scale public benchmark for the VI-ReID task, which contains 491 identities captured by four RGB cameras and two IR cameras in both outdoor and indoor environments. In this dataset, 22,258 RGB images and 11,909 IR images with 395 identities are collected for training. In the inference stage, the query set consists of 3,803 IR images with 96 identities and the galley set contains 301 randomly selected RGB images. RegDB is collected by an RGB camera and an IR camera, which contains 4,120 RGB images and 4,120 IR images with 412 identities. To be specific, the dataset is randomly divided into two non-overlapping sets: one set is used for training and the other is for testing.
Evaluation Protocols. The experiment follows the standard evaluation settings in VI-ReID, i.e., Cumulative Matching Characteristics (CMC)  Ye et al. (2018) and mean Average Precision (mAP). Specifically, we present the results of all-search and indoor-search modes for SYSU-MM01. The RegDB dataset contains two evaluation settings: Visible2Thermal and Thermal2Visible, i.e., Visible images as query images to search gallery sets consisting of infrared images and vice versa.
Implementation Details. The proposed method is implemented on four NVIDIA 3090 GPUs with PyTorch. We adopt the feature extractor in AGW Ye et al. (2022), which is initialized with ImageNet-pretrained weights to extract 2048-dimensional features. During the training stage, the input images are resized to 288$\times$144. We follow augmentations in  Ye et al. (2021) for data augmentation. In one batch, we randomly sample 16 person IDs, and each person ID samples 16 instances. We set $M$ to be 16 for computational convenience. We adopt Adam as an optimizer to train the model with an initial learning rate of 3.5e-4, decaying to 1/10 of the previous value times every 20 epochs. The number of epochs is 100, in which the first 50 epochs are trained by contrastive loss with the centroid prototype. For the last 50 epochs, we train the model by contrastive loss with both the hard and dynamic prototypes. $E_{CPCL}$ is 50. At the beginning of each epoch, we utilize the DBSCAN Ester et al. (1996) algorithm to generate pseudo labels. During the inference stage, we use the momentum backbone $\phi_{m}$ to extract features and take the features of the global average pooling layer to calculate cosine similarity for retrieval. The momentum value $\alpha$ and $\beta$ is set to 0.1 and 0.999, respectively. The temperature hyper-parameter $\tau$ is set to 0.05 and the weighting hyper-parameter $\lambda$ in Eq.(20) is 0.5.
@@ 4.2 Results and Analysis
To comprehensively evaluate our method, we compare our method with 20 supervised VI-ReID methods, 4 semi-supervised VI-ReID methods, and 9 unsupervised VI-ReID methods. The comparison results on the SYSU-MM01 and RegDB are reported in Tab. 1.
Comparison with USVI-ReID Methods. As shown in Tab. 1, our method achieves superior performance compared with state-of-the-art USVI-ReID methods. MMM Shi et al. (2024) is proposed to establish reliable cross-modality correspondences and is also the current best-performing method. Our method with MMM can achieve 65.9% in Rank-1 and 61.8% in mAP, which surpasses that of MMM by a large margin of 4.3% and 3.9%. Notably, our method even without MMM gains the best performance with 64.4% in Rank-1 and 58.7% in mAP. Although existing USVI-ReID methods mentioned in Tab. 1 have made great progress in the USVI-ReID task, the neglect of diverse individual information hinders their further improvement. They overlook diverse individual information, which often constitutes hard samples. Thus, we propose progressive contrastive learning with hard and dynamic prototypes to mine hard samples, which can guide the model to learn more robust and discriminative features.
Comparison with SSVI-ReID Methods. There are four SSVI-ReID methods proposed to alleviate the problem of labeling cost by using a part of data annotations and they have made some progress. However, the application of these methods in real scenarios is still limited due to the fact that data annotations are relatively expensive. Remarkably, our method’s performance without any annotations surpasses that of all methods relying on partial annotation.
Comparison with SVI-ReID Methods. Surprisingly, our method without annotation outperforms several SVI-ReID methods, e.g., DDAG Ye et al. (2020), AGW Ye et al. (2022), NFS Chen et al. (2021), LbA Park et al. (2021). This shows the immense competitiveness of PCLMP compared to SVI-ReID methods that rely on complete data labeling. The superior performance of PCLMP mainly benefits from the hard prototype and dynamic prototype contrastive learning. Additionally, it’s worth acknowledging that there is still a significant disparity between PCLMP and the state-of-the-art fully-supervised results.

### v2

@@ 3.4 Dynamic Prototype Contrastive Learning
Inspired by MoCo [49], we design dynamic prototype contrastive learning in order to preserve the intrinsic variety in sample features. DPCL comprises an online encoder $\phi_{0}$ and a momentum encoder $\phi_{m}$. The momentum encoder mirrors the structure of the online encoder, which is updated by the accumulated weights of the online encoder:
@@
where $\beta$ is a momentum coefficient that controls the update speed of the momentum encoder. $t$ and $t-1$ refer to the current and last iteration, respectively. The momentum encoder ${\phi}_{m}$ is updated by the moving averaged weights, which are resistant to sudden fluctuations or noisy updates [49].
We use the momentum encoder $\phi_{m}$ to extract feature representation and store them in visible memory $DM_{RGB}$={$dm^{v}_{1}$, $dm^{v}_{2}$, $\cdots$, $dm^{v}_{N_{v}}$} and infrared memory $DM_{IR}$={$dm^{r}_{1}$, $dm^{r}_{2}$, $\cdots$, $dm^{r}_{N_{r}}$}. We randomly select $M$ visible/infrared samples from each cluster, denoted as $X^{v}_{i}$ and $X^{r}_{j}$.as follows:
@@
$dm^{v}_{i}=\begin{cases}\underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\max}\left\|\phi_{m}(v_{j})-f^{v}_{i}\right\|&\text{if }y_{j}=y_{i}\\ \underset{{{\forall f^{v}_{i}\in F^{v}_{i}}}}{\arg\min}\left\|\phi_{m}(v_{j})-f^{v}_{i}\right\|&\text{if }y_{j}\neq y_{i}\end{cases},$ (19)
@@
$\mathcal{L}_{DPCL}^{v}=\frac{1}{N_{v}}\sum_{i\in N_{v}}\frac{-1}{|V_{i}|}\sum_{v\in V_{i}}\log{\frac{\text{exp}\left({\phi}_{m}(v)\cdot{dm}_{+}^{v}/\tau\right)}{\sum\limits_{j\in N_{v}}\text{exp}\left({{\phi}_{m}(v)}\cdot{dm}^{v}_{j}/\tau\right)}},$ (20)
$\mathcal{L}_{DPCL}^{r}=\frac{1}{N_{r}}\sum_{i\in N_{r}}\frac{-1}{|R_{i}|}\sum_{r\in R_{i}}\log{\frac{\text{exp}\left({\phi}_{m}(r)\cdot{dm}_{+}^{r}/\tau\right)}{\sum\limits_{j\in N_{r}}\text{exp}\left({{\phi}_{m}(r)}\cdot{dm}^{r}_{j}/\tau\right)}},$ (21)
$\mathcal{L}_{DPCL}=\mathcal{L}_{DPCL}^{v}+\mathcal{L}_{DPCL}^{r},$ (22)
@@ 3.5 Progressive Contrastive Learning
In the initial training phases, representations are generally of lower quality. Introducing hard samples at this period could be counterproductive, potentially leading the model optimization in an incorrect direction right from the start [40, 50]. To address this issue, we introduce the Progressive Contrastive Learning, which forms the overall loss function:
$\mathcal{L}_{PCLHD}=\left\{\begin{array}[]{l}\mathcal{L}_{CPCL},\quad\text{ if}\text{ epoch }\leqslant E_{\text{CPCL }}\\ \lambda\mathcal{L}_{HPCL}+(1-\lambda)\mathcal{L}_{DPCL},\quad\text{else}\end{array}\right.$ (23)
@@
4 Experiment
We conduct extensive experiments to validate the superiority of our proposed method. First, we provide the detailed experiment setting, which contains datasets, evaluation protocols, and implementation details. Then, we compare our method with many state-of-the-art VI-ReID methods and conduct ablation studies. In addition, to better illustrate our method, we also exhibit further analysis. If not specified, we conduct analysis experiments on SYSU-MM01 in the all-search mode.
@@ 4 Experiment
SVI-ReID DDAG [30] ECCV’20 54.8 53.0 61.0 68.0 69.4 63.5 68.1 61.8
AGW [51] TPAMI’21 47.5 47.7 54.2 63.0 70.1 66.4 70.5 65.9
NFS [52] CVPR’21 56.9 55.5 62.8 69.8 80.5 72.1 78.0 69.8
LbA [53] ICCV’21 55.4 54.1 58.5 66.3 74.2 67.6 72.4 65.5
CAJ [1] ICCV’21 69.9 66.9 76.3 80.4 85.0 79.1 84.8 77.8
MPANet [32] CVPR’21 70.6 68.2 76.7 81.0 83.7 80.9 82.8 80.7
DART [18] CVPR’22 68.7 66.3 72.5 78.2 83.6 75.7 82.0 73.8
FMCNet [29] CVPR’22 66.3 62.5 68.2 74.1 89.1 84.4 88.4 83.9
MAUM [28] CVPR’22 71.7 68.8 77.0 81.9 87.9 85.1 87.0 84.3
MID [54] AAAI’22 60.3 59.4 64.9 70.1 87.5 84.9 84.3 81.4
LUPI [55] ECCV’22 71.1 67.6 82.4 82.7 88.0 82.7 86.8 81.3
DEEN [56] CVPR’23 74.7 71.8 80.3 83.3 91.1 85.1 89.5 83.4
SGIEL [7] CVPR’23 77.1 72.3 82.1 83.0 92.2 86.6 91.1 85.2
PartMix [6] CVPR’23 77.8 74.6 81.5 84.4 85.7 82.3 84.9 82.5
CAL [57] ICCV’23 74.7 71.7 79.7 83.7 94.5 88.7 93.6 87.6
MUN [5] ICCV’23 76.2 73.8 79.4 82.1 95.2 87.2 91.9 85.0
SAAI [8] ICCV’23 75.9 77.0 83.2 88.0 91.1 91.5 92.1 92.0
FDNM [58] arXiv’24 77.8 75.1 87.3 89.1 95.5 90.0 94.0 88.7
PMWGCN [9] TIFS’24 90.6 84.5 88.8 81.6 66.8 64.9 72.6 76.2
LCNL [59] IJCV’24 70.2 68.0 76.2 80.3 85.6 78.7 84.0 76.9
SSVI-ReID OTLA [11] ECCV’22 48.2 43.9 47.4 56.8 49.9 41.8 49.6 42.8
TAA [13] TIP’23 48.8 42.3 50.1 56.0 62.2 56.0 63.8 56.5
DPIS [12] ICCV’23 58.4 55.6 63.0 70.0 62.3 53.2 61.5 52.7
USVI-ReID H2H [42] TIP’21 30.2 29.4 - - 23.8 18.9 - -
OTLA [11] ECCV’22 29.9 27.1 29.8 38.8 32.9 29.7 32.1 28.6
ADCA [14] MM’22 45.5 42.7 50.6 59.1 67.2 64.1 68.5 63.8
NGLR [60] MM’23 50.4 47.4 53.5 61.7 85.6 76.7 82.9 75.0
MBCCM [61] MM’23 53.1 48.2 55.2 62.0 83.8 77.9 82.8 76.7
CCLNet [45] MM’23 54.0 50.2 56.7 65.1 69.9 65.5 70.2 66.7
PGM [17] CVPR’23 57.3 51.8 56.2 62.7 69.5 65.4 69.9 65.2
GUR* [16] ICCV’23 61.0 57.0 64.2 69.5 73.9 70.2 75.0 69.9
MMM [48] arXiv’24 61.6 57.9 64.4 70.4 89.7 80.5 85.8 77.0
PCLHD Ours 64.4 58.7 69.5 74.4 84.3 80.7 82.7 78.4
PCLHD+MMM Enhanced 65.9 61.8 70.3 74.9 89.6 83.7 87.0 80.9
Table 2: Ablation studies on SYSU-MM01 in all search mode and indoor search mode. “Baseline” means the model trained following PGM  [17]. Rank-R accuracy(%) and mAP(%) are reported.
Component All Search Indoor Search
@@ 4.1 Experiment Setting
Dataset. We evaluate our method on two common benchmarks in VI-ReID: SYSU-MM01 [62] and RegDB [63]. SYSU-MM01 is a large-scale public benchmark for the VI-ReID task, which contains 491 identities captured by four RGB cameras and two IR cameras in both outdoor and indoor environments. In this dataset, 22,258 RGB images and 11,909 IR images with 395 identities are collected for training. In the inference stage, the query set consists of 3,803 IR images with 96 identities and the galley set contains 301 randomly selected RGB images. RegDB is collected by an RGB camera and an IR camera, which contains 4,120 RGB images and 4,120 IR images with 412 identities. To be specific, the dataset is randomly divided into two non-overlapping sets: one set is used for training and the other is for testing.
Evaluation Protocols. The experiment follows the standard evaluation settings in VI-ReID, i.e., Cumulative Matching Characteristics (CMC)  [64] and mean Average Precision (mAP).
Implementation Details.  We adopt the feature extractor in AGW [51], which is initialized with ImageNet-pretrained weights to extract 2048-dimensional features. During the training stage, the input images are resized to 288$\times$144. We follow augmentations in  [1] for data augmentation. In one batch, we randomly sample 16 person IDs, and each person ID samples 16 instances. We set $M$ to be 16 for computational convenience. The number of epochs is 100, in which the first 50 epochs are trained by contrastive loss with the centroid prototype. For the last 50 epochs, we train the model by contrastive loss with both the hard and dynamic prototypes. $E_{CPCL}$ is 50. At the beginning of each epoch, we utilize the DBSCAN [44] algorithm to generate pseudo labels. During the inference stage, we use the momentum encoder $\phi_{m}$ to extract features and take the features of the global average pooling layer to calculate cosine similarity for retrieval. The momentum value $\alpha$ and $\beta$ is set to 0.1 and 0.999, respectively. The temperature hyper-parameter $\tau$ is set to 0.05 and the weighting hyper-parameter $\lambda$ in Eq.(23) is 0.5.
@@ 4.2 Results and Analysis
To comprehensively evaluate our method, we compare our method with 18 supervised VI-ReID methods, 3 semi-supervised VI-ReID methods, and 9 unsupervised VI-ReID methods. The comparison results on the SYSU-MM01 and RegDB are reported in Tab. 1.
Comparison with USVI-ReID Methods. As shown in Tab. 1, our method achieves superior performance compared with state-of-the-art USVI-ReID methods. MMM [48] is proposed to establish reliable cross-modality correspondences and is also the current best-performing method. Our method with MMM can achieve 65.9% in Rank-1 and 61.8% in mAP, which surpasses that of MMM by a large margin of 4.3% and 3.9%. Notably, our method even without MMM gains the best performance with 64.4% in Rank-1 and 58.7% in mAP. Although existing USVI-ReID methods mentioned in Tab. 1 have made great progress in the USVI-ReID task, the neglects of divergence and variety hinder their further improvement. They overlook divergence and variety, which often constitutes hard samples. Thus, we propose progressive contrastive learning with hard and dynamic prototypes to mine hard samples, which can guide the model to learn more robust and discriminative features.
Comparison with SSVI-ReID Methods. There are three SSVI-ReID methods proposed to alleviate the problem of labeling cost by using a part of data annotations. Remarkably, our method’s performance without any annotations surpasses that of all methods relying on partial annotation.
Comparison with SVI-ReID Methods. Surprisingly, our method without annotation outperforms several SVI-ReID methods, e.g., DDAG [30], AGW [51], NFS [52], LbA [53]. This shows the immense competitiveness of PCLHD compared to SVI-ReID methods that rely on complete data labeling. The superior performance of PCLHD mainly benefits from the hard prototype and dynamic prototype contrastive learning. Additionally, it’s worth acknowledging that there is still a significant disparity between PCLHD and the state-of-the-art fully-supervised results.
4.3 Ablation Study
We conduct ablation studies on the SYSU-MM01 dataset in both all-search and indoor-search modes to show the effectiveness of each component in our method. The results are shown in Tab. 2.
Baseline Settings. We use PGM [17] as our baseline. Although PGM has achieved a promising performance on the USVI-ReID task, the neglect of hard samples hinders its further improvement.
Effectiveness of HPCL. The HPCL is proposed to mine divergence. As shown in Tab. 2, When adding the HPCL on Baseline, the performance improves a large margin of 5.8% in Rank-1 and 5.1% in mAP, respectively. It shows that divergence can be effectively mined using hard prototype contrast learning, facilitating the model to learn more discriminative features.
Effectiveness of DPCL. The DPCL is proposed to mine variety. The results show that Rank-1 accuracy can be improved by 2.8% in Rank-1 and 2.7% in mAP when adding the DPCL on Baseline, which confirms that contrastive learning with dynamic prototype can learn variety.
Effectiveness of PCL. PCL is introduced to smoothly shift the model’s attention from commonality to divergence and variety. The results show that Rank-1 accuracy can be improved by about 1% in Rank-1 and mAP compared to adding simultaneously the HPCL and DPCL on the Baseline. This confirms that progressive contrastive learning plays a valuable role in assisting HPCL and DPCL.
Surprisingly, contrastive learning with both hard and dynamic prototypes significantly exceeds the baseline by a large margin of 8.1% in Rank-1 and 7.0% in mAP. The HPCL and DPCL can complement each other to learn divergence and variety, which effectively guides the network to learn more robust and discriminative features.